OpenStudy (anonymous):

find a linear regression model this is a table Year|1980|'85|'90|'95|'96|'97|'98|'99 spending|4.5|6.6|11.6|12.6|12.7|13.1|12.9|12.4 in billions using x=0 as 1980, predict the spending for 2006

OpenStudy (anonymous):

and write and equation of the line of best fit (round to hundreths)

OpenStudy (dumbcow):

is this for a stats class?

OpenStudy (anonymous):

no its for geometry

OpenStudy (dumbcow):

well the easiest way is to find the line that goes through the 2 end points (0,4.5), (19,12.4) find slope and y_intercept

OpenStudy (anonymous):

where did the 19 come from?

OpenStudy (dumbcow):

year 1999 1999 - 1980 = 19

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

how does that predict the spending for 2006 though?

OpenStudy (dumbcow):

once you get the equation of the line find y when x=26 (2006-1980 = 26)

OpenStudy (anonymous):

ok so i got y=.42-4.5

OpenStudy (anonymous):

then i make x 26 and have y=.42(26)-4.5

OpenStudy (dumbcow):

correct good job however i wouldn't round the slope so much cause it will effect the answer in next part slope = .4158

OpenStudy (anonymous):

ok

OpenStudy (dumbcow):

oh wait oops what is the y_intercept? its the initial value right?

OpenStudy (anonymous):

the y intercept is 4.5

OpenStudy (dumbcow):

is it positive or negative 4.5

OpenStudy (anonymous):

positive

OpenStudy (dumbcow):

ok, you wrote "-4.5" just checking

OpenStudy (anonymous):

thanks

OpenStudy (dumbcow):

:)